# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v2

@@ Appendix A Appendix / supplemental material
A.2 Experiment Setup and Additional Result
A.2.1 MuJoCo Tasks
Data. Following the similar data generation pipeline in Brown et al. [2019], we generate the expert demonstrations and preference dataset as follows. We first train an expert agent by leveraging the ground-truth reward function and the popular Soft Actor-Critic (SAC) algorithm Haarnoja et al. [2018], which is developed to solve policy optimization problems with continuous action space. During the training process, we save the policy checkpoints and collect $30$k samples from each checkpoint. To achieve precise control of dataset quality, we categorize the data collected into three different classes: low-, medium-, and high-quality datasets according to the performance of the checkpoints. Then we combine the low- and medium-quality data as the preference dataset and use high-quality as demonstration data. Other details are given in Appendix A.2.
Settings. We set $L_{1}(\cdot)$ using the likelihood objective (9), $L_{2}(\cdot)$ as the BTL objective (11), and $L_{3}$ as KL-regularized objective given in (10). We implement the RLHF Brown et al. [2019] using the following 3 steps: (i) utilize the demonstration dataset to pre-train the policy using behavior cloning (BC) Pomerleau [1988]; (ii) conduct reward learning using preference data; (iii) implement the SAC for policy fine-tuning. In the proposed AIHF, the policy $\pi^{0}$ is initiated with a random policy, and the policy optimization problem is also solved using SAC.
In all Mujoco experiments, we assess the performance of benchmark algorithms on Hopper, HalfCheetah, and Walker2d from OpenAI Gym. To ensure a fair comparison, we utilize an open-source implementation of SAC [^5] https://github.com/openai/spinningup as the foundational RL algorithm for all methods. Regarding the choice of the penalty function $c(\cdot)$, since our proposed pipeline, AIHF , initiates with a random policy where $\pi(a|s)$ follows a uniform distribution, the kl penalty function is reduced to an entropy function.
The quality of the preference dataset and demonstration dataset are listed as follows Tab. 1:
Task Dataset Non-prefer Data Prefer Data Demonstration Data
Hopper-v2 $2345.20\pm 329.93$ $3024.63\pm 40.52$ $3559.61\pm 73.12$
Walker2d-v2 $3952.60\pm 444.45$ $5091.71\pm 291.73$ $5453.41\pm 71.07$
Table 1: The quality of preference and demonstration.
In the following sections, we include additional details about language alignment experiments.
A.2.2 Sentiment-Controlled Generation
Dataset Generation: In the IMDb sentiment completion task, we generate the demonstrations and preference datasets using the following procedure. Initially, we train a Language Model by employing the ground-truth reward function distilbert-imdb and the Proximal Policy Optimization (PPO) algorithm on 30% of the training dataset for IMDb. Throughout the training process, we save the policy checkpoint every 500 PPO steps. Subsequently, we select an additional 40% of the training dataset and generate a response for each prompt for each checkpoint. According to the evaluation score of each generation, we categorize the data collected into different classes: low-, medium-, and high-quality datasets, then we combine low-quality and medium-quality as preference datasets, and use high-quality as demonstration datasets. We show the average score of each dataset in 3,
[…]

### v3

@@ A.3.1 Useful Lemmas
A.3.3 Proof of Lemma A.1
To prove Lemma A.1, we will prove the equality (39a) and the equality (39b) respectively. The constants $L_{q}$ and $L_{c}$ in Lemma A.1 has the expression:
$L_{q}:=\frac{L_{r}}{1-\gamma},\quad L_{c}:=\frac{2L_{q}L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}+\frac{2L_{g}}{1-\gamma}.$
A.3.4 Proof of Inequality (39a)
In this subsection, we prove the inequality (39a) in Lemma A.1.
We show that $Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}$ has a bounded gradient with respect to any reward parameter $\theta$, then the inequality (39a) holds due to the mean value theorem. According to the soft Bellman equation, we have shown the explicit expression of $\nabla_{\theta}Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}(s,a)$ for any $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Using this expression, we have the following series of relations:
$\|\nabla_{\theta}Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}(s,a)\|$ $=\bigg{\|}\mathbb{E}_{a_{0}\sim\pi_{\theta}(\cdot|s_{0}),s_{1}\sim P(\cdot|s_{0},a_{0})}\Big{[}\nabla_{\theta}r(s_{0},a_{0};\theta)+\gamma\nabla_{\theta}V_{\theta}(s_{1})\Big{]}\bigg{\|}$
=$\overset{(i)}{=}\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)~{}\bigg{|}(s_{0},a_{0})=(s,a)\bigg{]}\bigg{\|}$
$\overset{(ii)}{\leq}\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\bigg{\|}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{\|}~{}\bigg{|}(s_{0},a_{0})=(s,a)\bigg{]}$
$\overset{(iii)}{\leq}\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}L_{r}~{}\bigg{|}(s_{0},a_{0})=(s,a)\bigg{]}$
$=\frac{L_{r}}{1-\gamma}$ (49)
where (i) is from the equality (47) in the proof of Lemma A.1, (ii) follows Jensen’s inequality and (iii) follows the inequality (38a) in Assumption 2. To complete this proof, we use the mean value theorem to show that
$|Q^{\text{soft}}_{r_{\theta_{1}},\pi_{\theta_{1}}}(s,a)-Q^{\text{soft}}_{r_{\theta_{2}},\pi_{\theta_{2}}}(s,a)|\leq\|\max_{\theta}\nabla_{\theta}Q^{\text{soft}}_{r_{\theta},\pi_{\theta}}(s,a)\|\cdot\|\theta_{1}-\theta_{2}\|\leq L_{q}\|\theta_{1}-\theta_{2}\|$ (50)
where the last inequality follows (49) and we denote $L_{q}:=\frac{L_{r}}{1-\gamma}$. Therefore, we have proved the Lipschitz continuous inequality in (39a).
A.3.5 Proof of Inequality (39b)
In this section, we prove the inequality (39b) in Lemma A.1.
According to Lemma A.1, the gradient $\nabla L_{1}(\theta)$ is expressed as:
$\nabla L_{1}(\theta)=\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta)\bigg{]}.$ (51)
Using the above relation, we have
$\|\nabla L_{1}(\theta_{1})-\nabla L_{1}(\theta_{2})\|$
=$\overset{(i)}{=}\bigg{\|}\bigg{(}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}\bigg{)}-$
$\quad\quad\bigg{(}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{\|}$
$\leq\underbrace{\bigg{\|}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}}_{\rm:=term~{}A}+$
$\quad\quad\underbrace{\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}}_{\rm:=term~{}B}$ (52)
where (i) follows the exact gradient expression in equation (51). Then we separately analyze term A and term B in (52).
For term A, it follows that
$\bigg{\|}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}$
$\overset{(i)}{\leq}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\big{\|}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})-\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\big{\|}\bigg{]}$
$\overset{(ii)}{\leq}\mathbb{E}_{\tau\sim\pi^{\rm E}}\bigg{[}\sum_{t\geq 0}\gamma^{t}L_{g}\|\theta_{1}-\theta_{2}\|\bigg{]}$
$=\frac{L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (53)
where (i) follows Jensen’s inequality and (ii) is from (38b) in Assumption 2.
For the term B, it holds that
$\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}$
$\overset{(i)}{\leq}\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{1}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}\bigg{\|}$
$\quad+\bigg{\|}\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{]}\bigg{\|}$
$\overset{(ii)}{\leq}\frac{1}{1-\gamma}\bigg{\|}\mathbb{E}_{(s,a)\sim d(\cdot,\cdot;\pi_{\theta_{1}})}\bigg{[}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}-\mathbb{E}_{(s,a)\sim d(\cdot,\cdot;\pi_{\theta_{2}})}\bigg{[}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{]}\bigg{\|}$
$\quad+\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{t\geq 0}\gamma^{t}\bigg{\|}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})-\nabla_{\theta}r(s_{t},a_{t};\theta_{2})\bigg{\|}\bigg{]}$
$\overset{(iii)}{\leq}\frac{1}{1-\gamma}\bigg{\|}\sum_{s\in\mathcal{S},a\in\mathcal{A}}\nabla_{\theta}r(s_{t},a_{t};\theta_{1})\bigg{(}d(s,a;\pi_{\theta_{1}})-d(s,a;\pi_{\theta_{2}})\bigg{)}\bigg{\|}+\mathbb{E}_{\tau\sim\pi_{\theta_{2}}}\bigg{[}\sum_{k\geq 0}\gamma^{k}L_{g}\|\theta_{1}-\theta_{2}\|\bigg{]}$
$\overset{(iv)}{\leq}\frac{2L_{r}}{1-\gamma}\|d(\cdot,\cdot;\pi_{\theta_{1}})-d(\cdot,\cdot;\pi_{\theta_{2}})\|_{TV}+\frac{L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (54)
where (i) follows the triangle inequality, (ii) is from Jensen’s inequality and the definition of the discounted state-action visitation measure $d(s,a;\pi):=(1-\gamma)\pi(a|s)\sum_{t\geq 0}\gamma^{t}\mathcal{P}^{\pi}(s_{t}=s|s_{0}\sim\eta)$; (iii) is from (38b) in Assumption 2;(iv) is from (38a) and the definition of the total variation norm.
Consider the $L_{2}$ term:
$L_{2}(\theta):=\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\left[\log\left(\sigma\left(R(\tau_{w};\theta)-R(\tau_{i};\theta)\right)\right)\right]$
where $\sigma(x)$ is sigmoid function defined by: $\sigma(x)=\frac{1}{1+e^{-x}}$. We have
$\nabla_{\theta}L_{2}(\theta)=\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\left[(1-\sigma(R(\tau_{w};\theta)-R(\tau_{i};\theta)))\cdot\left(\nabla_{\theta}R(\tau_{w};\theta)-\nabla_{\theta}R(\tau_{i};\theta)\right)\right].$
$=\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\big{[}\left(\nabla_{\theta}R(\tau_{w};\theta)-\nabla_{\theta}R(\tau_{i};\theta)\right)-\sigma(R(\tau_{w};\theta)-R(\tau_{i};\theta))(\nabla_{\theta}R(\tau_{w};\theta)-\nabla_{\theta}R(\tau_{i};\theta))\big{]}$
Using the triangle inequality, we obtain the following equation:
$\|\nabla L_{2}(\tau_{w},\tau_{l};\theta_{1})-\nabla L_{2}(\tau_{w},\tau_{l};\theta_{2})\|$
$\leq$ $\underbrace{\bigg{\|}\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\big{[}\left(\nabla_{\theta}R(\tau_{w};\theta_{1})-\nabla_{\theta}R(\tau_{i};\theta_{1})\right)-\left(\nabla_{\theta}R(\tau_{w};\theta_{2})-\nabla_{\theta}R(\tau_{i};\theta_{2})\right)\big{]}\bigg{\|}}_{\rm:=term~{}A}$
$+\underbrace{\bigg{\|}\mathbb{E}_{(\tau_{i},\tau_{w})\sim\pi^{P}}\big{[}\sigma(R(\tau_{w};\theta_{1})-R(\tau_{i};\theta_{1}))(\nabla_{\theta}R(\tau_{w};\theta_{1})-\nabla_{\theta}R(\tau_{i};\theta_{1}))}$
$\underbrace{-\sigma(R(\tau_{w};\theta_{2})-R(\tau_{i};\theta_{2}))(\nabla_{\theta}R(\tau_{w};\theta_{2})-\nabla_{\theta}R(\tau_{i};\theta_{2}))\big{]}\bigg{\|}}_{\rm:=term~{}B}$ (55)
First we bound the term A of (A.3.5)
termA=([$\rm term~{}A=\bigg{\|}\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{]}\bigg{)}-\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{\|}$
$\leq\bigg{\|}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})\bigg{\|}+\bigg{\|}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{\|}$
$\leq\frac{2L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (56)
Then we bounded term B of (A.3.5):
$\rm term~{}B=$
${=}\bigg{\|}\sigma\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{]}\bigg{)}\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{]}\bigg{)}$
$-\sigma\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{(}\bigg{[}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{]}\bigg{)}\bigg{\|}$
$=\bigg{\|}\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}$
$-\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}$
$+\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}$
$\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}\bigg{\|}$
$\leq\bigg{\|}\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})$
$\quad+\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{2})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{2})\bigg{)}\bigg{\|}+\bigg{\|}\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}$
$\quad\bigg{[}\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}-\sigma\bigg{(}\sum_{t\geq 0}\gamma^{t}\nabla_{\theta}r(s^{w}_{t},a^{w}_{t};\theta_{1})-\gamma^{t}\nabla_{\theta}r(s^{l}_{t},a^{l}_{t};\theta_{1})\bigg{)}\bigg{]}\bigg{\|}$
$\leq\frac{2L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|+\frac{L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|=\frac{3L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$ (57)
Plugging the inequalities (53), (54) to (52), it holds that
$\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\|$
$\leq\frac{2L_{r}}{1-\gamma}\|d(\cdot,\cdot;\pi_{\theta_{1}})-d(\cdot,\cdot;\pi_{\theta_{2}})\|_{TV}+\frac{6L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$
$\overset{(i)}{\leq}\frac{2L_{r}C_{d}}{1-\gamma}\|Q^{\text{soft}}_{r_{\theta_{1}},\pi_{\theta_{1}}}-Q^{\text{soft}}_{r_{\theta_{2}},\pi_{\theta_{2}}}\|+\frac{6L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$
$\overset{(ii)}{\leq}\frac{2L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}\|Q^{\text{soft}}_{r_{\theta_{1}},\pi_{\theta_{1}}}-Q^{\text{soft}}_{r_{\theta_{2}},\pi_{\theta_{2}}}\|_{\infty}+\frac{6L_{g}}{1-\gamma}\|\theta_{1}-\theta_{2}\|$
$\overset{(iii)}{\leq}\bigg{(}\frac{2L_{q}L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}+\frac{6L_{g}}{1-\gamma}\bigg{)}\|\theta_{1}-\theta_{2}\|.$ (58)
Define the constant $L_{c}:=\frac{2L_{q}L_{r}C_{d}\sqrt{|\mathcal{S}|\cdot|\mathcal{A}|}}{1-\gamma}+\frac{5L_{g}}{1-\gamma}$, we have the following inequality:
$\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\|\leq L_{c}\|\theta_{1}-\theta_{2}\|.$
Therefore, we complete the proof of the inequality (39b) in Lemma A.1.
A.4 Proof of Theorem 4.1
In this section, we prove (20a) and (20b) respectively, to show the convergence of the lower-level problem and the upper-level problem.
A.4.1 Proof of (20a)
In this proof, we first show the convergence of the lower-level variable $\{\pi_{k}\}_{k\geq 0}$. Recall that we approximate the optimal policy $\pi_{\theta_{k}}$ by $\pi_{k+1}$ at each iteration $k$. We first analyze the approximation error between $\pi_{\theta_{k}}$ and $\pi_{k+1}$ as follows. For any $s\in\mathcal{S}$ and $a\in\mathcal{A}$, we have the following relation:
$\big{|}\log\big{(}\pi_{k+1}(a|s)\big{)}-\log\big{(}\pi_{\theta_{k}}(a|s)\big{)}\big{|}$
=$\overset{(i)}{=}\bigg{|}\log\bigg{(}\frac{\pi^{0}(a|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)\big{)}}{\sum_{\tilde{a}}\exp\pi^{0}(\tilde{a}|s)\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})\big{)}}\bigg{)}-\log\bigg{(}\frac{\pi^{0}(a|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)\big{)}}{\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})\big{)}}\bigg{)}\bigg{|}$
$\overset{(ii)}{\leq}\big{|}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)\big{|}+\bigg{|}\log\bigg{(}\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})\big{)}\bigg{)}$
$\quad-\log\bigg{(}\sum_{\tilde{a}}\pi^{0}(\tilde{a}|s)\exp\big{(}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})\big{)}\bigg{)}\bigg{|}$ (59)
where (i) follows (40); (ii) is by the triangle inequality. We further analyze the second term in (59).
We first denote the operator $\log(\|w\exp(v)\|_{1}):=\log(\|\sum_{\tilde{a}\in\mathcal{A}}w\exp(v_{\tilde{a}})\|_{1})$, where the vector $w,v\in\mathbb{R}^{|\mathcal{A}|}$ and $v=[v_{1},v_{2},\cdots,v_{|\mathcal{A}|}],w=[w_{1},w_{2},\cdots,w_{|\mathcal{A}|}]$. Then for any $v^{\prime},v^{\prime\prime}\in\mathbb{R}^{|\mathcal{A}|}$, we have the following relation:
$\big{|}\log\big{(}\|w^{\prime}\exp(v^{\prime})\|_{1}\big{)}-\log\big{(}\|w^{\prime\prime}\exp(v^{\prime\prime})\|_{1}\big{)}$ =$\overset{(i)}{=}\big{\langle}v^{\prime}-v^{\prime\prime},\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}|_{v=v^{c}}\big{\rangle}$
$\leq\|v^{\prime}-v^{\prime\prime}\|_{\infty}\cdot\|\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}|_{v=v^{c}}\|_{1}$
=$\overset{(ii)}{=}\|v^{\prime}-v^{\prime\prime}\|_{\infty}$ (60)
where (i) follows the mean value theorem and $v_{c}$ is a convex combination of $v^{\prime}$ and $v^{\prime\prime}$; (ii) follows the following equalities:
$[\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}]_{i}=\frac{w_{i}\exp(v_{i})}{\sum_{1\leq a\leq|\mathcal{A}|}w_{a}\exp(v_{a})},\quad\|\nabla_{v}\log\big{(}\|w\exp(v)\|_{1}\big{)}\|_{1}=1,\quad\forall v\in\mathbb{R}^{|\mathcal{A}|}.$
Through plugging (60) into (59), it holds that
$\big{|}\log\big{(}\pi_{k+1}(a|s)\big{)}-\log\big{(}\pi_{\theta_{k}}(a|s)\big{)}\big{|}$
$\leq\big{|}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)\big{|}+\max_{\tilde{a}\in\mathcal{A}}\big{|}Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,\tilde{a})-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,\tilde{a})\big{|}$ (61)
Taking the infinity norm over $\mathbb{R}^{|\mathcal{S}|\cdot|\mathcal{A}|}$, the following result holds:
$\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}\leq 2\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$ (62)
where $\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}=\max_{s\in\mathcal{S},a\in\mathcal{A}}|\log\pi_{k+1}(a|s)-\log\pi_{\theta_{k}}(a|s)|$ and $\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}=\max_{s\in\mathcal{S},a\in\mathcal{A}}|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}(s,a)-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}(s,a)|$.
Based on the inequality (62), we analyze $\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$ to show the convergence of the policy estimates. It leads to the following analysis:
$\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{k}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}\|_{\infty}$
$=\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}+Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}+Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}\|_{\infty}$
$\leq\|Q^{\text{soft}}_{r_{\theta_{k}},\pi_{\theta_{k}}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}\|_{\infty}$
$\overset{(i)}{\leq}L_{q}\|\theta_{k}-\theta_{k-1}\|+\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+\|Q_{r_{\theta_{k}},\pi_{k}}^{\text{soft}}-Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}\|_{\infty}$
$\overset{(ii)}{\leq}\|Q_{r_{\theta_{k-1}},\pi_{k}}^{\text{soft}}-Q^{\text{soft}}_{r_{\theta_{k-1}},\pi_{\theta_{k-1}}}\|_{\infty}+2L_{q}\|\theta_{k}-\theta_{k-1}\|$ (63)
where (i) is from (39a) in Lemma A.1; (ii) follows (39a). Based on (63), we further analyze the two terms in (63) as below.
Recall we have the “soft” Bellman operator expressed as below:
[…]